Two balls are drawn in succession without replacement from a box containing 4 red balls and 6 blue balls. Let Z be the random variable representing the number of blue balls. Construct the probability distribution of the random variable Z.
Answer by Edwin McCravy(20086)

First determine the probabilities of the events.

  events           Probability
    RR          (4/10)(3/9) = 2/15
    RB          (4/10)(6/9) = 4/15
    BR          (6/10)(4/9) = 4/15 
    BB          (6/10)(5/9) = 1/3

The probability of 0 blue balls (RR)is 2/15
The probability of 1 blue ball is (RB or BR) is 4/15+4/15 = 8/15
The probability of 2 blue balls (BB) is 1/3

So the probability distribution is:

Z   p(Z)
--------   
0   2/15
1   8/15
2   1/3

Notice that the sum of the probabilities = 2/15+8/15+1/3 = 1

Edwin